OpenStudy (anonymous):

Which is the vertex of the graph of y=x2+5x+6 Answers: A. 2.5,0.25 B.2,3 C.-3,-2 D.-2.5,-0.25

hero (hero):

The vertex occurs at the intersection of the line of symmetry and the graph of y.

OpenStudy (anonymous):

...

hero (hero):

@Caylahill, where does the line of symmetry occur?

OpenStudy (anonymous):

take the derivative of y=x2+5x+6 2x+5=0 =>x=-5/2=-2.5 y=(-2.5)^2+5(-2.5)+6 y=6.25-12.5+6=-0.25 (-2.5,-0.25)

hero (hero):

@AJ01, good answer, except @Caylahill probably has no idea what a derivative is.